OpenStudy (anonymous):

Help! What is the polynomial function of lowest degree with lead coefficient 1 and roots i, –2, and 2? f(x)=x^3-x^2-4x+4 f(x)=x^4-3x^2-4 f(x)=x^4+3x^2-4 f(x)=x^3+x^2-4x-4

OpenStudy (danjs):

Well , for the first one... x=2 or x=-2 or x=1 is that i supposed to be a one?

OpenStudy (anonymous):

No, it is supposed to be i, the imaginary number

OpenStudy (danjs):

well it has 3 roots, so it is a 3rd degree polynomial

OpenStudy (anonymous):

so already A or D?

OpenStudy (danjs):

The first one has roots, 2 , -2, 1 The Last one has roots 2 , -2 , -1

OpenStudy (danjs):

oh i see, i wrote it wrong

OpenStudy (danjs):

The second one has the roots: 2, -2, +i , -i

OpenStudy (anonymous):

The last one or the second one?

OpenStudy (danjs):

1) x=2, -2, 1 2) x= 2, -2, +i, -i 3) x = +1, -1, +2i, -2i 4) x = 2, -2, -1

OpenStudy (anonymous):

so b is the correct answer

OpenStudy (danjs):

yea